OpenStudy (anonymous):

-3b - 5 + 4 greater than or equal to -4

OpenStudy (anonymous):

-3b - 5 + 4 > or = -4 -3b - 1 > or = -4 -3b > or = -3 b > or = 1

OpenStudy (anonymous):

wait im sorry its b < or = 1 you flip the symbol when you divide or multiply by a negative number sorry

OpenStudy (anonymous):

so b is less than or equal to 1